A cuboidal hole of cross-section area 10cm square is cut out of a larger cube of side x cm.
Show that the volume of the remaining solid can be expressed as v=x square (x square-10)
Given that the volume this solid is 800 cm square, find x, giving your answer to 2 dp.

Answer by mananth(16946)

A cuboidal hole of cross-section area 10cm square is cut out of a larger cube of side x cm.
Show that the volume of the remaining solid can be expressed as v=x square (x square-10)
Given that the volume this solid is 800 cm square, find x, giving your answer to 2 dp.
..
volume of cuboidal hole = 10x cm^3
..
volume of cuboid = x^3
difference = x^3-10x
difference = x(x^2-10)
...
Volume of the cuboidal portion is 800 cm^3
...
10x= 800x= 80 the height. so also the lengthe of the edge of the cube.